# LLM for Barcodes: Generating Diverse Synthetic Data for Identity Documents

### v1

@@ Introduction
A barcode is a machine readable representation of data in the form of parallel lines or patterns. It is used to store information about the object to which it’s attached and play a crucial role in how we manage and process information today. Barcodes make it possible to store, retrieve, and verify data quickly and accurately, which is essential in many industries. There are several types of barcodes, with the two primary categories being one-dimensional (1D) and two-dimensional (2D) barcodes. Barcodes play a crucial role in identity documents like insurance cards, Driver’s license, and ID cards. They store essential personal information in a machine readable format, facilitating quick and accurate verification. These documents are issued by governments or authorized organizations to confirm someone’s identity. Barcodes in these documents help improve security, make verification faster, and allow different systems to work together more efficiently. However, building reliable models to detect barcodes in identity documents is a tough problem. One major challenge is the lack of large and diverse datasets needed to train these models. Privacy laws, like the General Data Protection Regulation (GDPR), restrict access to real-world identity documents, and because these documents often contain sensitive personal information, there are additional restrictions. Furthermore, identity documents vary a lot in design as they can have different data encoded, layouts, languages, and barcode types depending on the country or organization that issued them. This makes it hard for models trained on limited data to work well across all types of documents.
Some existing tools attempt to generate synthetic data to bridge this gap, but they come with limitations. For instance, tools like Faker rely on predefined fields, which don’t adapt well to the variability and complexity of real-world identity documents. This approach often results in data that lacks diversity and fails to accurately represent actual documents. For example, driver’s licenses from states like New Jersey and Texas encode different types of information, and the structure may vary not only between documents but also within a single document. As a result, models trained on such synthetic data may struggle to perform effectively. To solve these problems, we propose a new way to create synthetic datasets using LLMs. These models can generate realistic and diverse data for identity documents without relying on predefined fields. The data reflects the variety of global identity documents, including differences in culture and region, while also ensuring that privacy is protected. This data is then converted into barcodes and overlayed onto document templates, creating a dataset that looks and feels real but doesn’t include any sensitive personal information.
@@
Large Language Models are playing an increasingly important role in synthetic data generation, helping to address data scarcity and improve machine learning applications. For instance, (Long et al. 2024) provides a detailed overview of LLM-driven workflows for creating synthetic data, along with the challenges they face. (Patel, Raffel, and Callison-Burch 2024) introduces DataDreamer, a tool designed to ensure that synthetic data generation using LLMs is reproducible. Similarly, (Gupta et al. 2024) showcases TarGEN, which uses targeted prompts to enhance the quality of generated datasets. On a larger scale, (Ge et al. 2024) presents Persona Hub, a framework for creating synthetic personas at scale. Lastly, (Veselovsky et al. 2023) explores methods for generating reliable synthetic data for social science research using grounding techniques. Together, these works demonstrate the power of LLMs in producing diverse, high-quality synthetic data for a wide range of purposes
@@
This section outlines how data is generated for various types of identity documents, including Driver’s Licenses, Insurance Cards, and University IDs. Each document type is designed to include specific information tailored to its purpose. For example, Driver’s Licenses typically feature details like the holder’s name, address, license number, date of birth, and issuing state, though the exact fields can vary between states. Insurance Cards, on the other hand, include information such as the policy number, coverage dates, provider, and plan type, which often differ based on the insurance company. Similarly, University IDs contain data like the student’s name, ID number, department, and enrollment year, with variations depending on the institution. These differences highlight the need for flexible and realistic metadata generation to reflect the diversity of real-world documents. Our approach consists of three key components: data generation using a LLM and the Faker library, barcode encoding with the pyBarcode library, and integrating these barcodes into document templates. To further enrich the dataset, we apply data augmentation techniques, ensuring greater diversity and realism in the generated documents.

### v2

@@ Introduction
A barcode is a machine readable representation of data in the form of parallel lines or patterns. It is used to store information about the object to which it’s attached and play a crucial role in how we manage and process information today. Barcodes make it possible to store, retrieve, and verify data quickly and accurately, which is essential in many industries. There are several types of barcodes, with the two primary categories being one-dimensional (1D) and two-dimensional (2D) barcodes. Barcodes play a crucial role in identity documents like insurance cards, Driver’s license, and ID cards. They store essential personal information in a machine readable format, facilitating quick and accurate verification. These documents are issued by governments or authorized organizations to confirm someone’s identity. Barcodes in these documents help improve security, make verification faster, and allow different systems to work together more efficiently. However, building reliable models to detect barcodes in identity documents is a tough problem. One major challenge is the lack of large and diverse datasets needed to train these models. Privacy laws, like the General Data Protection Regulation (GDPR), restrict access to real-world identity documents, and because these documents often contain sensitive personal information, there are additional restrictions. Furthermore, identity documents vary a lot in design as they can have different data encoded, layouts, languages, and barcode types depending on the country or organization that issued them. This makes it hard for models trained on limited data to work well across all types of documents (Agarwal et al. 2024b).
Some existing tools attempt to generate synthetic data to bridge this gap (Agarwal, Panda, and Pachauri 2024; Agarwal et al. 2024a), but they come with limitations. For instance, tools like Faker rely on predefined fields, which don’t adapt well to the variability and complexity of real-world identity documents. This approach often results in data that lacks diversity and fails to accurately represent actual documents. For example, driver’s licenses from states like New Jersey and Texas encode different types of information, and the structure may vary not only between documents but also within a single document. As a result, models trained on such synthetic data may struggle to perform effectively. To solve these problems, we propose a new way to create synthetic datasets using LLMs. These models can generate realistic and diverse data for identity documents without relying on predefined fields. The data reflects the variety of global identity documents, including differences in culture and region, while also ensuring that privacy is protected. This data is then converted into barcodes and overlayed onto document templates, creating a dataset that looks and feels real but doesn’t include any sensitive personal information.
@@
Large Language Models are playing an increasingly important role in synthetic data generation, helping to address data scarcity and improve machine learning applications (Agarwal et al. 2024c). For instance, (Long et al. 2024) provides a detailed overview of LLM-driven workflows for creating synthetic data, along with the challenges they face. (Patel, Raffel, and Callison-Burch 2024) introduces DataDreamer, a tool designed to ensure that synthetic data generation using LLMs is reproducible. Similarly, (Gupta et al. 2024) showcases TarGEN, which uses targeted prompts to enhance the quality of generated datasets. On a larger scale, (Ge et al. 2024) presents Persona Hub, a framework for creating synthetic personas at scale. Lastly, (Veselovsky et al. 2023) explores methods for generating reliable synthetic data for social science research using grounding techniques. Together, these works demonstrate the power of LLMs in producing diverse, high-quality synthetic data for a wide range of purposes
@@
This section outlines how data is generated for various types of identity documents, including Driver’s Licenses, Insurance Cards, and University IDs. Each document type is designed to include specific information tailored to its purpose. For example, Driver’s Licenses typically feature details like the holder’s name, address, license number, date of birth, and issuing state, though the exact fields can vary between states. Insurance Cards, on the other hand, include information such as the policy number, coverage dates, provider, and plan type, which often differ based on the insurance company. Similarly, University IDs contain data like the student’s name, ID number, department, and enrollment year, with variations depending on the institution. These differences highlight the need for flexible and realistic metadata generation to reflect the diversity of real-world documents. Our approach consists of three key components: data generation using a LLM and the Faker library, barcode encoding with the pyBarcode library, and integrating these barcodes into document templates. To further enrich the dataset, we apply data augmentation techniques (Agarwal et al. 2024a), ensuring greater diversity and realism in the generated documents.